SW&IT Meeting Minutes 2018-07-16
Attendees: Andrew, David, Doug, Igor, Kellee, Matthieu
, Petr, Stephen, Xianyu, Yang
News from management
Previous Action Items
Bring any topics you want for discussion at the meeting here. Use <topic>[<name>] formatting.
This week in github
- Working on IT issue #7325 to develop a tool for mining the collimation data.
- After collimation data mining tool, study per instrument collimation with Petr (based on Doug's hypothesis).
- Finished working with Doug on transitioning pointing model improvements, learned a lot of information.
- Working on pointing model related wiki pages.
- Continue to study the AO supervisor code and the IT issue #7261 (memory problem with WfsArbitrator)
- programed the bootloader through ICD3 programmer;
- loaded the processor program to the processor;
- Read the memory map by Matlab and sent comments to MG;
- PBP plan discussion with Guido;
- LBTI SOUL network and Jumbo frame dicussion;
- helped AGW1 packing;
- Disabled DX AdSec motoring scripts;
- Prepared the SX AdSec interface for the potential 4D work;
- Supported the DX AdSec lab test.
- Write the PBP requirement document with Guido;
- Identify the PBP functionalities with Guido;
- some other summer shutdown activities.
- DNS configuration in VirtualBox
- LBTI soul network discussion.
- Firewall configuration.
- Rearrangement of CRA to accomodate SOUL, re-rack and removed some excess inventory.
- racked SOUL computer for tests on Thursday/Friday.
- research and quote for SOUL switch.
- replaced SSD card in Synology
- Dataprobe configuration look-see.
- Ilya allsky.
- prepared OVMS SD card, ready to install in the ractangle
- learned finish command for gdb (missed that for ages)
- rediscovered /\t in vim to check for tabs, and retab command
- TTL pull request, problems with MODS (Luci seems to work reasonable well, but to my surprise the TPK library does not count with changes in rotator speed as telescope approach local meridian)
- collimation-data tool (with Yang)
- Reviewed Petr's update to the SysLog code in TCS and wrote up some notes for future development.
- OVMS kernel update discussions with Petr to arrange a plan and time. Rebuilt the OVMS real-time code with the updated kernel and put it on the RACK for the test.
- git meeting #2 to discuss workflow.
- Final GCS checkout with the new OAC in Tucson. Pushed my updates to a TCS branch in git.
- Wrote up some notes for Stephen to check out an LBC dataprobe on the mountain.
- Vacation on Thursday.
- Support update to the OVMS kernel - vacation moved to next week.
- Take instruments out of Alarm Handler and Status Server for the summer, other shutdown conveniences.
- Vacation on Thursday this week through Tuesday next week.
I'll be in on Thursday next week, 26-July.
This week has been LUCI
In particular working on issuetraks:
6944: AFC daemon service freezes.
7323: Adding anamorphic magnification for LUCI1 (DONE)
7329: AFC Shifting groups failure
This is an aladin issue, we know where the problem is but unsure as to what or how to solve it.
7330: Media Clipping added to AFC routine
This was added
7331: AFC Exceeds max allowed shift
- Worked on MODS laptop drawing into Vault project
- Convoluted; must import files through several versions of software to modernize them
- Will set up intern this week so he can experience tedium
- Progress on IT email lists
- Instruments needs some extra confirmation but is largely ready to go
- Chasing managers
- Install Centos VM for Petr (done)
- MODS drawings
- Vacation Thursday & Friday
- Test scripts for swing arm script
- Close out handover items with MPIA
- Final draft of handover document
- Run simulator from docker containers
- Configure swing arm script to run on jet machine